CLISP-READLINE(3) CLISP-READLINE(3) NAME readline−getalinefromauserwithediting NOTATION Anemacs-stylenotationisusedtodenotekeystrokes.ControlkeysaredenotedbyC−key,e.g., C−nmeansControl−N.Similarly,metakeysaredenotedbyM−key,soM−xmeansMeta−X. (Onkeyboardswithoutametakey,M−xmeansESCx,i.e.,presstheEscapekeythenthexkey. ThismakesESCthemetaprefix.ThecombinationM−C−xmeansESC−Control−x,orpressthe EscapekeythenholdtheControlkeywhilepressingthexkey.) Readlinecommandsmaybegivennumericarguments,whichnormallyactasarepeatcount. Sometimes,however,itisthesignoftheargumentthatissignificant.Passinganegativeargumenttoacommandthatactsintheforwarddirection(e.g.,kill−line)causesthatcommandto actinabackwarddirection.Commandswhosebehaviorwithargumentsdeviatesfromthisare noted. Whenacommandisdescribedaskillingtext,thetextdeletedissavedforpossiblefutureretrieval (yanking).Thekilledtextissavedinakillring.Consecutivekillscausethetexttobeaccumulatedintooneunit,whichcanbeyankedallatonce.Commandswhichdonotkilltextseparate thechunksoftextonthekillring. INITIALIZATIONFILE Readlineiscustomizedbyputtingcommandsinaninitializationfile(theinputrcfile).Thename ofthisfileistakenfromthevalueoftheINPUTRCenvironmentvariable.Ifthatvariableis unset,thedefaultis˜ /.inputrc.Whenaprogramwhichusesthereadlinelibrarystartsup,the initfileisread,andthekeybindingsandvariablesareset.Thereareonlyafewbasicconstructs allowedinthereadlineinitfile.Blanklinesareignored.Linesbeginningwitha#arecomments. Linesbeginningwitha$indicateconditionalconstructs.Otherlinesdenotekeybindingsand variablesettings.Eachprogramusingthislibrarymayadditsowncommandsandbindings. Forexample,placing M−Control−u:universal−argument or C−Meta−u:universal−argument intotheinputrcwouldmakeM−C−uexecutethereadlinecommanduniversal−argument. Thefollowingsymboliccharacternamesarerecognizedwhileprocessingkeybindings:RUBOUT, DEL,ESC,LFD,NEWLINE,RET,RETURN,SPC,SPACE,andTAB. Inadditiontocommandnames,readlineallowskeystobeboundtoastringthatisinsertedwhen thekeyispressed(amacro). KeyBindings Thesyntaxforcontrollingkeybindingsintheinputrcfileissimple.Allthatisrequiredisthe nameofthecommandorthetextofamacroandakeysequencetowhichitshouldbebound. Thenamemaybespecifiedinoneoftwoways:asasymbolickeyname,possiblywithMeta−or Control−prefixes,orasakeysequence.Whenusingtheformkeyname:function-nameormacro, keynameisthenameofakeyspelledoutinEnglish.Forexample: Control−u:universal−argument Meta−Rubout:backward−kill−word Control−o:">&output" Intheaboveexample,C−uisboundtothefunctionuniversal−argument,M-DELisboundto thefunctionbackward−kill−word,andC−oisboundtorunthemacroexpressedontheright handside(thatis,toinsertthetext>&outputintotheline). Inthesecondform,"keyseq":function−nameormacro,keyseqdiffersfromkeynameabovein thatstringsdenotinganentirekeysequencemaybespecifiedbyplacingthesequencewithindoublequotes.SomeGNUEmacsstylekeyescapescanbeused,asinthefollowingexample. GNU 2000January23 1 CLISP-READLINE(3) CLISP-READLINE(3) "\C−u":universal−argument "\C−x\C−r":re−read−init−file "\e[11˜ ":"FunctionKey1" Inthisexample,C-uisagainboundtothefunctionuniversal−argument.C-xC-risboundto thefunctionre−read−init−file,andESC[11˜ isboundtoinsertthetextFunctionKey1. ThefullsetofGNUEmacsstyleescapesequencesis \C−controlprefix \M−metaprefix \eanescapecharacter \\backslash \"literal" \’literal’ InadditiontotheGNUEmacsstyleescapesequences,asecondsetofbackslashescapesisavailable: \aalert(bell) \bbackspace \ddelete \fformfeed \nnewline \rcarriagereturn \thorizontaltab \vverticaltab \nnnthecharacterwhoseASCIIcodeistheoctalvaluennn(onetothreedigits) \xnnnthecharacterwhoseASCIIcodeisthehexadecimalvaluennn(onetothreedigits) Whenenteringthetextofamacro,singleordoublequotesshouldbeusedtoindicateamacro definition.Unquotedtextisassumedtobeafunctionname.Inthemacrobody,thebackslash escapesdescribedaboveareexpanded.Backslashwillquoteanyothercharacterinthemacro text,including"and’. Bashallowsthecurrentreadlinekeybindingstobedisplayedormodifiedwiththebindbuiltin command.Theeditingmodemaybeswitchedduringinteractiveusebyusingthe−ooptionto thesetbuiltincommand.Otherprogramsusingthislibraryprovidesimilarmechanisms.The inputrcfilemaybeeditedandre-readifaprogramdoesnotprovideanyothermeanstoincorpo-ratenewbindings. Variables Readlinehasvariablesthatcanbeusedtofurthercustomizeitsbehavior.Avariablemaybeset intheinputrcfilewithastatementoftheform setvariable−namevalue Exceptwherenoted,readlinevariablescantakethevaluesOnorOff.Thevariablesandtheir defaultvaluesare: bell−style(audible) Controlswhathappenswhenreadlinewantstoringtheterminalbell.Ifsettonone, readlineneverringsthebell.Ifsettovisible,readlineusesavisiblebellifoneisavailable.Ifsettoaudible,readlineattemptstoringtheterminal’sbell. comment−begin(‘‘;’’) Thestringthatisinsertedinvimodewhentheinsert−commentcommandisexecuted. ThiscommandisboundtoM−#inemacsmodeandto#invicommandmode. completion−ignore−case(Off) IfsettoOn,readlineperformsfilenamematchingandcompletioninacase−insensitive fashion. GNU 2000January23 2 CLISP-READLINE(3) CLISP-READLINE(3) blink−matching−paren(On) IfsettoOn,readlinewilldisplaythecorrespondingopeningparenthesiswhenaclosing parenthesisistyped. completion−query−items(100) Thisdetermineswhentheuserisqueriedaboutviewingthenumberofpossiblecompletionsgeneratedbythepossible−completionscommand.Itmaybesettoanyinteger valuegreaterthanorequaltozero.Ifthenumberofpossiblecompletionsisgreaterthan orequaltothevalueofthisvariable,theuserisaskedwhetherornothewishestoview them;otherwisetheyaresimplylistedontheterminal. disable−completion(Off) IfsettoOn,readlinewillinhibitwordcompletion.Completioncharacterswillbe insertedintothelineasiftheyhadbeenmappedtoself-insert. editing−mode(emacs) Controlswhetherreadlinebeginswithasetofkeybindingssimilartoemacsorvi.editing−modecanbesettoeitheremacsorvi. enable−keypad(Off) WhensettoOn,readlinewilltrytoenabletheapplicationkeypadwhenitiscalled. Somesystemsneedthistoenablethearrowkeys. expand−tilde(Off) Ifsettoon,tildeexpansionisperformedwhenreadlineattemptswordcompletion. horizontal−scroll−mode(Off) WhensettoOn,makesreadlineuseasinglelinefordisplay,scrollingtheinputhorizontallyonasinglescreenlinewhenitbecomeslongerthanthescreenwidthratherthan wrappingtoanewline. keymap(emacs) Setthecurrentreadlinekeymap.Thesetoflegalkeymapnamesisemacs,emacs-stan-dard,emacs-meta,emacs-ctlx,vi,vi-move,vi-command,andvi-insert.viisequivalentto vi-command;emacsisequivalenttoemacs-standard.Thedefaultvalueisemacs;the valueofediting−modealsoaffectsthedefaultkeymap. mark−directories(On) IfsettoOn,completeddirectorynameshaveaslashappended. mark−modified−lines(Off) IfsettoOn,historylinesthathavebeenmodifiedaredisplayedwithaprecedingasterisk (*). convert−meta(On) IfsettoOn,readlinewillconvertcharacterswiththeeighthbitsettoanASCIIkey sequencebystrippingtheeighthbitandprependinganescapecharacter(ineffect,using escapeasthemetaprefix). meta−flag(Off) IfsettoOn,readlinewillenableeight-bitinput(thatis,itwillnotstripthehighbit fromthecharactersitreads),regardlessofwhattheterminalclaimsitcansupport. output−meta(Off) IfsettoOn,readlinewilldisplaycharacterswiththeeighthbitsetdirectlyratherthanas ameta-prefixedescapesequence. print−completions−horizontally(Off) IfsettoOn,readlinewilldisplaycompletionswithmatchessortedhorizontallyinalphabeticalorder,ratherthandownthescreen. show−all−if−ambiguous(Off) Thisaltersthedefaultbehaviorofthecompletionfunctions.Ifsettoon,wordswhich havemorethanonepossiblecompletioncausethematchestobelistedimmediately insteadofringingthebell. visible−stats(Off) IfsettoOn,acharacterdenotingafile’stypeasreportedbystat(2)isappendedtothe filenamewhenlistingpossiblecompletions. GNU 2000January23 3 CLISP-READLINE(3) CLISP-READLINE(3) ConditionalConstructs ReadlineimplementsafacilitysimilarinspirittotheconditionalcompilationfeaturesoftheC preprocessorwhichallowskeybindingsandvariablesettingstobeperformedastheresultoftests. Therearefourparserdirectivesused. $ifThe$ifconstructallowsbindingstobemadebasedontheeditingmode,theterminal beingused,ortheapplicationusingreadline.Thetextofthetestextendstotheendof theline;nocharactersarerequiredtoisolateit. modeThemode=formofthe$ifdirectiveisusedtotestwhetherreadlineisinemacs orvimode.Thismaybeusedinconjunctionwiththesetkeymapcommand, forinstance,tosetbindingsintheemacs-standardandemacs-ctlxkeymapsonlyif readlineisstartingoutinemacsmode. termTheterm=formmaybeusedtoincludeterminal-specifickeybindings,perhaps tobindthekeysequencesoutputbytheterminal’sfunctionkeys.Thewordon therightsideofthe=istestedagainstthefullnameoftheterminalandthe portionoftheterminalnamebeforethefirst−.Thisallowssuntomatchboth sunandsun−cmd,forinstance. application Theapplicationconstructisusedtoincludeapplication-specificsettings.Each programusingthereadlinelibrarysetstheapplicationname,andaninitialization filecantestforaparticularvalue.Thiscouldbeusedtobindkeysequencesto functionsusefulforaspecificprogram.Forinstance,thefollowingcommandadds akeysequencethatquotesthecurrentorpreviouswordinBash: $ifbash #Quotethecurrentorpreviousword "\C-xq":"\eb\"\ef\"" $endif $endifThiscommand,asseeninthepreviousexample,terminatesan$ifcommand. $elseCommandsinthisbranchofthe$ifdirectiveareexecutedifthetestfails. $include Thisdirectivetakesasinglefilenameasanargumentandreadscommandsandbindings fromthatfile.Forexample,thefollowingdirectivewouldread/etc/inputrc: $include/etc/inputrc SEARCHING Readlineprovidescommandsforsearchingthroughthecommandhistoryforlinescontaininga specifiedstring.Therearetwosearchmodes:incrementalandnon-incremental. Incrementalsearchesbeginbeforetheuserhasfinishedtypingthesearchstring.Aseachcharacterofthesearchstringistyped,readlinedisplaysthenextentryfromthehistorymatchingthe stringtypedsofar.Anincrementalsearchrequiresonlyasmanycharactersasneededtofindthe desiredhistoryentry.TheEscapecharacterisusedtoterminateanincrementalsearch.Control-Jwillalsoterminatethesearch.Control-Gwillabortanincrementalsearchandrestoretheorigi-nalline.Whenthesearchisterminated,thehistoryentrycontainingthesearchstringbecomes thecurrentline.Tofindothermatchingentriesinthehistorylist,typeControl-SorControl-Ras appropriate.Thiswillsearchbackwardorforwardinthehistoryforthenextlinematchingthe searchstringtypedsofar.Anyotherkeysequenceboundtoareadlinecommandwillterminate thesearchandexecutethatcommand.Forinstance,anewlinewillterminatethesearchand accepttheline,therebyexecutingthecommandfromthehistorylist. Non-incrementalsearchesreadtheentiresearchstringbeforestartingtosearchformatchinghis-torylines.Thesearchstringmaybetypedbytheuserorbepartofthecontentsofthecurrent line. GNU 2000January23 4 CLISP-READLINE(3) CLISP-READLINE(3) EDITINGCOMMANDS Thefollowingisalistofthenamesofthecommandsandthedefaultkeysequencestowhichthey arebound.Commandnameswithoutanaccompanyingkeysequenceareunboundbydefault. CommandsforMoving beginning−of−line(C−a) Movetothestartofthecurrentline. end−of−line(C−e) Movetotheendoftheline. forward−char(C−f) Moveforwardacharacter. backward−char(C−b) Movebackacharacter. forward−word(M−f) Moveforwardtotheendofthenextword.Wordsarecomposedofalphanumericcharacters(lettersanddigits). backward−word(M−b) Movebacktothestartofthecurrentorpreviousword.Wordsarecomposedofalphanumericcharacters(lettersanddigits). clear−screen(C−l) Clearthescreenleavingthecurrentlineatthetopofthescreen.Withanargument, refreshthecurrentlinewithoutclearingthescreen. redraw−current−line Refreshthecurrentline. CommandsforManipulatingtheHistory accept−line(Newline,Return) Acceptthelineregardlessofwherethecursoris.Ifthislineisnon-empty,addittothe historylist.Ifthelineisamodifiedhistoryline,thenrestorethehistorylinetoitsoriginalstate. previous−history(C−p) Fetchthepreviouscommandfromthehistorylist,movingbackinthelist. next−history(C−n) Fetchthenextcommandfromthehistorylist,movingforwardinthelist. beginning−of−history(M−<) Movetothefirstlineinthehistory. end−of−history(M−>) Movetotheendoftheinputhistory,i.e.,thelinecurrentlybeingentered. reverse−search−history(C−r) Searchbackwardstartingatthecurrentlineandmoving‘up’throughthehistoryasnecessary.Thisisanincrementalsearch. forward−search−history(C−s) Searchforwardstartingatthecurrentlineandmoving‘down’throughthehistoryasnecessary.Thisisanincrementalsearch. non−incremental−reverse−search−history(M−p) Searchbackwardthroughthehistorystartingatthecurrentlineusinganon-incremental searchforastringsuppliedbytheuser. non−incremental−forward−search−history(M−n) Searchforwardthroughthehistoryusinganon-incrementalsearchforastringsupplied bytheuser. history−search−forward Searchforwardthroughthehistoryforthestringofcharactersbetweenthestartofthe currentlineandthecurrentcursorposition(thepoint).Thisisanon-incrementalsearch. history−search−backward Searchbackwardthroughthehistoryforthestringofcharactersbetweenthestartofthe currentlineandthepoint.Thisisanon-incrementalsearch. GNU 2000January23 5 CLISP-READLINE(3) CLISP-READLINE(3) yank−nth−arg(M−C−y) Insertthefirstargumenttothepreviouscommand(usuallythesecondwordonthepreviousline)atpoint(thecurrentcursorposition).Withanargumentn,insertthenthword fromthepreviouscommand(thewordsinthepreviouscommandbeginwithword0).A negativeargumentinsertsthenthwordfromtheendofthepreviouscommand. yank−last−arg(M−.,M−) Insertthelastargumenttothepreviouscommand(thelastwordoftheprevioushistory entry).Withanargument,behaveexactlylikeyank−nth−arg.Successivecallsto yank−last−argmovebackthroughthehistorylist,insertingthelastargumentofeach lineinturn. CommandsforChangingText delete−char(C−d) Deletethecharacterunderthecursor.Ifpointisatthebeginningoftheline,thereare nocharactersintheline,andthelastcharactertypedwasnotboundtodelete−char, thenreturnEOF. backward−delete−char(Rubout) Deletethecharacterbehindthecursor.Whengivenanumericargument,savethedeleted textonthekillring. quoted−insert(C−q,C−v) Addthenextcharacterthatyoutypetothelineverbatim.ThisishowtoinsertcharacterslikeC−q,forexample. tab−insert(M-TAB) Insertatabcharacter. self−insert(a,b,A,1,!,...) Insertthecharactertyped. transpose−chars(C−t) Dragthecharacterbeforepointforwardoverthecharacteratpoint.Pointmovesforward aswell.Ifpointisattheendoftheline,thentransposethetwocharactersbeforepoint. Negativeargumentsdon’twork. transpose−words(M−t) Dragthewordbehindthecursorpastthewordinfrontofthecursormovingthecursor overthatwordaswell. upcase−word(M−u) Uppercasethecurrent(orfollowing)word.Withanegativeargument,uppercasethepreviousword,butdonotmovepoint. downcase−word(M−l) Lowercasethecurrent(orfollowing)word.Withanegativeargument,lowercasethepreviousword,butdonotmovepoint. capitalize−word(M−c) Capitalizethecurrent(orfollowing)word.Withanegativeargument,capitalizethepreviousword,butdonotmovepoint. KillingandYanking kill−line(C−k) Killthetextfromthecurrentcursorpositiontotheendoftheline. backward−kill−line(C−xRubout) Killbackwardtothebeginningoftheline. unix−line−discard(C−u) Killbackwardfrompointtothebeginningoftheline.Thekilledtextissavedonthekillring. kill−whole−line Killallcharactersonthecurrentline,nomatterwherethecursoris. kill−word(M−d) Killfromthecursortotheendofthecurrentword,orifbetweenwords,totheendofthe nextword.Wordboundariesarethesameasthoseusedbyforward−word. GNU 2000January23 6 CLISP-READLINE(3) CLISP-READLINE(3) backward−kill−word(M−Rubout) Killthewordbehindthecursor.Wordboundariesarethesameasthoseusedbybackward−word. unix−word−rubout(C−w) Killthewordbehindthecursor,usingwhitespaceasawordboundary.Thewordboundariesaredifferentfrombackward−kill−word. delete−horizontal−space(M−\) Deleteallspacesandtabsaroundpoint. kill−region Killthetextbetweenthepointandmark(savedcursorposition).Thistextisreferredto astheregion. copy−region−as−kill Copythetextintheregiontothekillbuffer. copy−backward−word Copythewordbeforepointtothekillbuffer.Thewordboundariesarethesameas backward−word. copy−forward−word Copythewordfollowingpointtothekillbuffer.Thewordboundariesarethesameas forward−word. yank(C−y) Yankthetopofthekillringintothebufferatthecursor. yank−pop(M−y) Rotatethekillring,andyankthenewtop.Onlyworksfollowingyankoryank−pop. NumericArguments digit−argument(M−0,M−1,...,M−−) Addthisdigittotheargumentalreadyaccumulating,orstartanewargument.M−− startsanegativeargument. universal−argument Thisisanotherwaytospecifyanargument.Ifthiscommandisfollowedbyoneormore digits,optionallywithaleadingminussign,thosedigitsdefinetheargument.Ifthecommandisfollowedbydigits,executinguniversal−argumentagainendsthenumericargument,butisotherwiseignored.Asaspecialcase,ifthiscommandisimmediatelyfollowedbyacharacterthatisneitheradigitorminussign,theargumentcountforthenext commandismultipliedbyfour.Theargumentcountisinitiallyone,soexecutingthis functionthefirsttimemakestheargumentcountfour,asecondtimemakestheargument countsixteen,andsoon. Completing complete(TAB) Attempttoperformcompletiononthetextbeforepoint.Theactualcompletionperformedisapplication-specific.Clisp,forinstance,attemptscompletionofsymbolnames, offunctionnames(ifthetextisprefixedby(or#’)oroffilenames(ifthetextisprefixed by#").Bash,ontheotherhand,attemptscompletiontreatingthetextasavariable(if thetextbeginswith$),username(ifthetextbeginswith˜ ),hostname(ifthetextbegins with@),orcommand(includingaliasesandfunctions)inturn.Ifnoneoftheseproduces amatch,filenamecompletionisattempted.Gdb,finally,allowscompletionofprogram functionsandvariables,andonlyattemptsfilenamecompletionundercertaincircumstances. possible−completions(M−?) Listthepossiblecompletionsofthetextbeforepoint. insert−completions(M−*) Insertallcompletionsofthetextbeforepointthatwouldhavebeengeneratedbypossible−completions. GNU 2000January23 7 CLISP-READLINE(3) CLISP-READLINE(3) menu−complete Similartocomplete,butreplacesthewordtobecompletedwithasinglematchfromthe listofpossiblecompletions.Repeatedexecutionofmenu−completestepsthroughthe listofpossiblecompletions,insertingeachmatchinturn.Attheendofthelistofcompletions,thebellisrungandtheoriginaltextisrestored.Anargumentofnmovesn positionsforwardinthelistofmatches;anegativeargumentmaybeusedtomovebackwardthroughthelist.ThiscommandisintendedtobeboundtoTAB,butisunbound bydefault. KeyboardMacros start−kbd−macro(C−x() Beginsavingthecharacterstypedintothecurrentkeyboardmacro. end−kbd−macro(C−x)) Stopsavingthecharacterstypedintothecurrentkeyboardmacroandstorethedefinition. call−last−kbd−macro(C−xe) Re-executethelastkeyboardmacrodefined,bymakingthecharactersinthemacro appearasiftypedatthekeyboard. Miscellaneous re−read−init−file(C−xC−r) Readinthecontentsoftheinputrcfile,andincorporateanybindingsorvariableassignmentsfoundthere. abort(C−g) Abortthecurrenteditingcommandandringtheterminal’sbell(subjecttothesettingof bell−style). do−uppercase−version(M−a,M−b,M−x,...) Ifthemetafiedcharacterxislowercase,runthecommandthatisboundtothecorrespondinguppercasecharacter. prefix−meta(ESC) Metafythenextcharactertyped.ESCfisequivalenttoMeta−f. undo(C−,C−xC−u) Incrementalundo,separatelyrememberedforeachline. revert−line(M−r) Undoallchangesmadetothisline.Thisislikeexecutingtheundocommandenough timestoreturnthelinetoitsinitialstate. tilde−expand(M−˜ ) Performtildeexpansiononthecurrentword. set−mark(C−@,M-<space>) Setthemarktothecurrentpoint.Ifanumericargumentissupplied,themarkissetto thatposition. exchange−point−and−mark(C−xC−x) Swapthepointwiththemark.Thecurrentcursorpositionissettothesavedposition, andtheoldcursorpositionissavedasthemark. character−search(C−]) Acharacterisreadandpointismovedtothenextoccurrenceofthatcharacter.Anegativecountsearchesforpreviousoccurrences. character−search−backward(M−C−]) Acharacterisreadandpointismovedtothepreviousoccurrenceofthatcharacter.A negativecountsearchesforsubsequentoccurrences. insert−comment(M−#) Thevalueofthereadlinecomment−beginvariableisinsertedatthebeginningofthe currentline,andthelineisacceptedasifanewlinehadbeentyped.Thismakesthecurrentlineashellcomment. GNU 2000January23 8 CLISP-READLINE(3) CLISP-READLINE(3) dump−functions Printallofthefunctionsandtheirkeybindingstothereadlineoutputstream.Ifa numericargumentissupplied,theoutputisformattedinsuchawaythatitcanbemade partofaninputrcfile. dump−variables Printallofthesettablevariablesandtheirvaluestothereadlineoutputstream.Ifa numericargumentissupplied,theoutputisformattedinsuchawaythatitcanbemade partofaninputrcfile. dump−macros Printallofthereadlinekeysequencesboundtomacrosandthestringstheyouput.Ifa numericargumentissupplied,theoutputisformattedinsuchawaythatitcanbemade partofaninputrcfile. emacs−editing−mode(C−e) Wheninvieditingmode,thiscausesaswitchtoemacseditingmode. vi−editing−mode(M−C−j) Wheninemacseditingmode,thiscausesaswitchtovieditingmode. DEFAULTKEYBINDINGS Thefollowingisalistofthedefaultemacsandvibindings.Characterswiththe8thbitsetare writtenasM−<character>,andarereferredtoasmetafiedcharacters.TheprintableASCIIcharactersnotmentionedinthelistofemacsstandardbindingsareboundtotheself−insertfunction, whichjustinsertsthegivencharacterintotheinputline.Inviinsertionmode,allcharactersnot specificallymentionedareboundtoself−insert.Charactersassignedtosignalgenerationby stty(1)ortheterminaldriver,suchasC-ZorC-C,retainthatfunction.Upperandlowercase metafiedcharactersareboundtothesamefunctionintheemacsmodemetakeymap.The remainingcharactersareunbound,whichcausesreadlinetoringthebell(subjecttothesettingof thebell−stylevariable). EmacsMode EmacsStandardbindings "C-@" set-mark "C-A" beginning-of-line "C-B" backward-char "C-D" delete-char "C-E" end-of-line "C-F" forward-char "C-G" abort "C-H" backward-delete-char "C-I" complete "C-J" accept-line "C-K" kill-line "C-L" clear-screen "C-M" accept-line "C-N" next-history "C-P" previous-history "C-Q" quoted-insert "C-R" reverse-search-history "C-S" forward-search-history "C-T" transpose-chars "C-U" unix-line-discard "C-V" quoted-insert "C-W" unix-word-rubout "C-Y" yank "C-]" character-search "C-" undo GNU 2000January23 9 CLISP-READLINE(3) CLISP-READLINE(3) ""to"/" self-insert "0"to"9" self-insert ":"to"˜ " self-insert "C-?" backward-delete-char EmacsMetabindings "M-C-G" abort "M-C-H" backward-kill-word "M-C-I" tab-insert "M-C-J" vi-editing-mode "M-C-M" vi-editing-mode "M-C-R" revert-line "M-C-Y" yank-nth-arg "M-C-[" complete "M-C-]" character-search-backward "M-space" set-mark "M-#" insert-comment "M-&" tilde-expand "M-*" insert-completions "M--" digit-argument "M-." yank-last-arg "M-0" digit-argument "M-1" digit-argument "M-2" digit-argument "M-3" digit-argument "M-4" digit-argument "M-5" digit-argument "M-6" digit-argument "M-7" digit-argument "M-8" digit-argument "M-9" digit-argument "M-<" beginning-of-history "M-=" possible-completions "M->" end-of-history "M-?" possible-completions "M-B" backward-word "M-C" capitalize-word "M-D" kill-word "M-F" forward-word "M-L" downcase-word "M-N" non-incremental-forward-search-history "M-P" non-incremental-reverse-search-history "M-R" revert-line "M-T" transpose-words "M-U" upcase-word "M-Y" yank-pop "M-\" delete-horizontal-space "M-˜ " tilde-expand "M-C-?" backward-delete-word "M-" yank-last-arg EmacsControl-Xbindings "C-XC-G" abort "C-XC-R" re-read-init-file GNU 2000January23 10 CLISP-READLINE(3) CLISP-READLINE(3) "C-XC-U" undo "C-XC-X" exchange-point-and-mark "C-X(" start-kbd-macro "C-X)" end-kbd-macro "C-XE" call-last-kbd-macro "C-XC-?" backward-kill-line VIModebindings VIInsertModefunctions "C-D" vi-eof-maybe "C-H" backward-delete-char "C-I" complete "C-J" accept-line "C-M" accept-line "C-R" reverse-search-history "C-S" forward-search-history "C-T" transpose-chars "C-U" unix-line-discard "C-V" quoted-insert "C-W" unix-word-rubout "C-Y" yank "C-[" vi-movement-mode "C-" undo ""to"˜ " self-insert "C-?" backward-delete-char VICommandModefunctions "C-D" vi-eof-maybe "C-E" emacs-editing-mode "C-G" abort "C-H" backward-char "C-J" accept-line "C-K" kill-line "C-L" clear-screen "C-M" accept-line "C-N" next-history "C-P" previous-history "C-Q" quoted-insert "C-R" reverse-search-history "C-S" forward-search-history "C-T" transpose-chars "C-U" unix-line-discard "C-V" quoted-insert "C-W" unix-word-rubout "C-Y" yank "" forward-char "#" insert-comment "$" end-of-line "%" vi-match "&" vi-tilde-expand "*" vi-complete "+" next-history "," vi-char-search GNU 2000January23 11 CLISP-READLINE(3) CLISP-READLINE(3) "-" previous-history "." vi-redo "/" vi-search "0" beginning-of-line "1"to"9" vi-arg-digit ";" vi-char-search "=" vi-complete "?" vi-search "A" vi-append-eol "B" vi-prev-word "C" vi-change-to "D" vi-delete-to "E" vi-end-word "F" vi-char-search "G" vi-fetch-history "I" vi-insert-beg "N" vi-search-again "P" vi-put "R" vi-replace "S" vi-subst "T" vi-char-search "U" revert-line "W" vi-next-word "X" backward-delete-char "Y" vi-yank-to "\" vi-complete "ˆ " vi-first-print "" vi-yank-arg "‘" vi-goto-mark "a" vi-append-mode "b" vi-prev-word "c" vi-change-to "d" vi-delete-to "e" vi-end-word "f" vi-char-search "h" backward-char "i" vi-insertion-mode "j" next-history "k" prev-history "l" forward-char "m" vi-set-mark "n" vi-search-again "p" vi-put "r" vi-change-char "s" vi-subst "t" vi-char-search "u" undo "w" vi-next-word "x" vi-delete "y" vi-yank-to "|" vi-column "˜ " vi-change-case GNU 2000January23 12 CLISP-READLINE(3) CLISP-READLINE(3) SEEALSO TheGnuReadlineLibrary,BrianFox TheGnuHistoryLibrary,BrianFox bash(1) FILES ˜ /.inputrc Individualreadlineinitializationfile AUTHORS BrianFox,FreeSoftwareFoundation bfox@gnu.org ChetRamey,CaseWesternReserveUniversity chet@ins.CWRU.Edu BrunoHaible haible@clisp.cons.org BUGREPORTS Ifyoufindabuginreadline,youshouldreportit.Sendmailtohaible@clisp.cons.org. BUGS It’stoobigandtooslow. GNU 2000January23 13